OpenStudy (anonymous):

What is the equation of the line that passes through the point (-2, 4)?

OpenStudy (anonymous):

I got y=-2, but I'm not sure if that's right.

jimthompson5910 (jim_thompson5910):

You're just given one point?

OpenStudy (phi):

You need 2 points to define a unique line. If you have one point, there are lots (an infinite number) of lines that you can draw through the point 

OpenStudy (phi):

But if you want just some line through (-2,4) you can pick the line x=-2 (notice -2 is the x value in (-2,4) ) or y=4 there are many other lines (equations) we could pick, but those are the easiest to write down.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I think I got it now. :)

OpenStudy (phi):

If you wanted a line with a slope of 1 that goes through (-2,4) you would use this formula \[ y - y_0= m(x-x0) \] where \( (x_0,y_0) \) is a point on the line with m=1, x0= -2 and y0= 4 we have y -4 =1(x- (-2)) y -4 = x +2 y = x+6 that is another line.